Bob's friend Tim is 4 times as old as he is. Six years from now, Tim will be twice as old as Bob will be. How old is Bob now? Algebra 1.
Answer by oscargut(686)

Let Bob's age = x
And Tim's age = y
then y=4x because (Tim is 4 times as old as Bob is)
y+6=2(x+6)(Six years from now, Tim will be twice as old as Bob will be)
then
4x+6=2x+12
then
2x=6
x=3
and y = 12
Answer: Bob is 3 years now